# Interpretable Fine-Tuning and Error Indication for Graph Neural Network Surrogate Models

### v3

@@ 3.2.3 Optimal Selection of Regularization Factor
The discussion above reveals the key role played by the regularization factor $\lambda$ as a controller for the trade-off between error tagging capability (quantitatively measured by the mean-squared error budget $\cal B$) and predictive accuracy of the fine-tuned GNN. The presence of such a trade-off – a direct by-product of the form of the regularized objective function in Eq. 5 – naturally leads to the notion of whether or not an ideal (or optimal) $\lambda$ that balances accuracy with error-tagging capability can be determined both a-priori and in a problem-independent manner. In particular, the following question comes into play: for a given Top-K reduction factor, is it possible to determine a value for $\lambda$ that maximizes the error budget contained in masked field (i.e., maximizes error-tagging capability) while also retaining the predictive capability of the baseline GNN? A $\lambda$-selection strategy to this end is described below.
To guide selection of $\lambda$, the expression in Eq. 5 can be simplified under two limiting conditions, both of which represent idealized performance targets of the regularized fine-tuned model. The first condition coincides with maximizing the MSE budget: in an ideal error-tagging scenario, 100% of the error in the forecast should be contained in the nodes identified by the masked field (a budget $\cal B$ of 1). In the ${\cal B}\rightarrow 1$ limit, the budget regularization loss ${\cal L}_{\cal B}$ also becomes 1. The second condition operates under the assumption of zero error in the fine-tuned model; in other words, an ideal model in the predictive sense would also recover ${\cal L}_{R}=0$ (see Eq. 5). Under these two limiting conditions, a simplified expression for the modeling objective in Eq. 5 can be obtained as $0={\cal L}_{\text{MSE}}+\lambda$, which in-turn provides an estimation for an ”ideal” $\lambda$ as $\lambda=-L_{\text{MSE}}$. Since a baseline model is available, the value $L_{\text{MSE}}$ can be set to $L_{\text{BL}}$, the converged MSE loss of the trained baseline GNN.
Although a negative value for $\lambda$ ($\lambda=-L_{\text{BL}}$) appears counter-intuitive, the relationship stems directly from the ideal modeling limits described above. Since $\lambda$ is negative, the budget regularization term during fine-tuning is adjusted to the direct budget instead of the inverse budget (i.e., ${\cal L}_{B}={\langle{\cal B}\rangle}$ instead of ${\cal L}_{B}=\frac{1}{\langle{\cal B}\rangle}$) to penalize large negative values for $\lambda{\cal L}_{B}$ during training. Although not shown here, the training loss histories for the MSE contribution (${\cal L}_{\text{MSE}}$, the single-step forecast error) validate this approach, in that the corresponding fine-tuned model in the $\lambda=-L_{\text{BL}}$ setting was found match the baseline prediction error on the training set.
@@ 3.3 Masked Fields on Extrapolated Geometries
The results in Sec. 3.1 show how the fine-tuning procedure provides the GNN the ability to generate identifiable structures through the masked fields. The results in Sec. 3.2 build on this intuition: through a budget regularization procedure conditioned on the identified masked fields, it was shown how a baseline GNN can be explicitly fine-tuned for interpretable error-tagging. Although these capabilities have many useful modeling implications (alongside the direct feature extraction benefits of the masked fields), trends in the previous sections were assessed only from the perspective of a single geometry, namely the backward-facing step.
@@
The objective of this section is to expand on the previous analyses by means of geometry extrapolation, which serves to not only verify masked fields trends, but also highlights the principal advantage of fully graph-based architectures: their ability to be used on unseen meshes. Here, fine-tuned GNNs trained on the BFS geometry are evaluated on two additional unseen geometries exhibiting different flow separation physics. These new geometries are the angled ramp and mounted cube configurations, which constitute a more challenging extrapolation task than, for example, extrapolating on another BFS geometry with the step positioned in a different part of the domain (the latter does not change the mechanism of flow separation, but rather only its spatial location). Fig. 14 in Appendix A provides images of the meshes used for the new geometries. Within the ramp and cube extrapolation context, the goal is to assess consistency in the masked fields from two angles, described below.
@@
Feature Identification and Error Tagging: To highlight error correlations, Fig. 11 provides instantaneous masked fields and corresponding spatial distributions of velocity component errors for the three aforementioned geometries: the backward-facing step, the ramp, and the wall-mounted cube. Analogous to Fig. 10, the figure includes outputs from fine-tuned GNNs trained with different budget regularization factors on the BFS geometry.
@@
For the BFS configuration in Fig. 11(a), the fine-tuned GNNs identify coherent structures downstream of the step. A similar identification of coherent structures is seen for the new geometries, in similar parts of the domain relative to the geometric feature. Most notably, the shift in the masked field towards the step anchor with increasing $\lambda$ – observed originally in BFS configuration – is also observed in the ramp configuration in Fig. 11(b). This shift is not seen in the cube case, which shows a markedly lower sensitivity to $\lambda$ than the other two geometries. Despite this, the GNN appears to be identifying coherent structures even in the extrapolated geometries, with the spatial location of these structures generally correlating with high levels of forecasting error. Additionally, the same concentration effect provided by the budget regularization strategy (which is the increase in the amount of error contained in the masked fields with $\lambda$) observed in the BFS case is observed in the ramp geometry – this effect is again much less apparent in the cube geometry. Although the masked fields are still picking up features in the mounted cube configuration, the lack of $\lambda$-sensitivity in Fig. 11(c) is likely due to the high absolute errors observed in the respective forecasts. Unlike the BFS and ramp cases, the cube-induced flowfield is not only governed by separation, but also the physical effects of flow impingement upstream of the geometry. Additionally, as described in Appendix A, the range in cells resolutions observed in the mesh for the cube configuration is greater than the other two cases, which may also be contributing to the higher errors. Ultimately, however, the extrapolations show how the masked fields largely preserve their feature identification capability in unseen geometries and largely remain correlated with regions of spatially concentrated forecasting error.
@@
Mask Decoherence During Rollout: Alongside analysis of spatial distributions and error tagging capability, it is of particular interest to evaluate whether the trends in the fine-tuned GNNs during rollout also translate to extrapolated geometries. More specifically, recall that through Fig. 7(d), presence of high-frequency noisy artifacts in the masked fields were shown to be intrinsically tied to the emergence of rollout errors in the BFS configuration. This emergence of noise is referred to as ”decoherency”. If a similar decoherency is observed during rollout for different geometries, the relationship between such a coherency breakdown and the presence of prediction errors is further verified (i.e., it strengthens confidence in the utility the masked fields provide in terms of predicting the error itself).
@@ 4 Conclusion
This work introduces an interpretable fine-tuning strategy for graph neural network based surrogate models, with focus on unstructured grid based fluid dynamics. More specifically, given a baseline GNN, the fine-tuning strategy consists of (a) appending an interpretable module to the GNN architecture through a skip connection, and (b) an isolated training stage to optimize the parameters introduced by this appended module. The resultant fine-tuned GNN adds interpretability to the baseline GNN by means of a graph pooling layer that acts as an adaptive sub-graph sampler, where these sub-graphs are identified in accordance to the underlying objective function (which here is the mesh-based forecasting error). Through this pooling operation, it is shown how the GNN forward pass generates masked fields as a by-product, which isolates regions in physical space in which these sampled sub-graphs are active. The features contained in these masked fields allow one to interpret the relationship between the input flow field at the current time instant, and the model-derived forecast at a future time instant. Through visualization of the masked fields, the fine-tuned GNNs allow one to access regions in physical space intrinsically linked to the modeling task (forecasting), thereby making the model behavior directly accessible from the perspective of the objective function used during training. As a result, the structures identified by pooling operation serve as an accessible link between the forecast-based objective function, the baseline model architecture, and the underlying physics of the problem being modeled.
@@
Additionally, by adding a regularization term to the objective function during the fine-tuning procedure in the form of an error budget, it was shown how the mask fields can also be used to identify, during inference, nodes that describe a majority of the anticipated forecasting error. Since the masked fields are an explicit function of the input flowfield (and not the predicted future flowfield), this regularization step adds a form of interpretable error-tagging capability to the modeling framework. A strategy to guide optimal selection of the regularization hyperparameter was also introduced, resulting in a novel approach that adds error-tagging capability to baseline GNNs without sacrificing baseline prediction accuracy. Additional analysis was performed on unseen geometries, with the goal of verifying the identification and error-tagging capability of the masked fields in challenging extrapolation scenarios. The analysis revealed how key GNN properties – particularly error localization and mask field decoherency – can indeed translate to unseen geometries.
@@
The authors believe that the fine-tuning strategies introduced here begin to address the limitations of existing black-box mesh-based modeling strategies, and can readily be extended to other applications both within and outside of fluid dynamics surrogate modeling. For example, the objective functions studied here concentrated on single-step forecasting errors – an interesting direction for future work is to append the same interpretable fine-tuning module, but include more rollout steps during the fine-tuning process, thereby producing an interpretable stability enhancement. Including physics-based loss terms is another promising direction, as it would allow one to investigate the relationship between the identified structures and governing equation residuals. Additionally, although the modeling context in this work focused on a node-level regression task, extension of the ideas presented here for graph-level (i.e., more macroscopic) regression tasks may be possible. Regarding treatment of multiple geometries, a promising direction is to fine-tune a GNN for interpretability on geometry B using a baseline trained on geometry A. These aspects, as well as other directions, will be explored in future work.
@@ 5 Acknowledgements
This research used resources of the Argonne Leadership Computing Facility, which is a U.S. Department of Energy Office of Science User Facility operated under contract DE-AC02-06CH11357. SB acknowledges funding support from the AETS fellowship from Argonne National Laboratory, provided by the Director, Office of Science, of the U.S. Department of Energy under contract DE-AC02-06CH11357. RM acknowledges funding support from ASCR for DOE-FOA-2493 “Data-intensive scientific machine learning” and from the RAPIDS2 Scientific Discovery through Advanced Computing (SciDAC) program.
@@ Appendix C Multiscale Message Passing Layer
As illustrated in Fig. 1, the baseline GNN architecture and the fine-tuning module leverage multiscale message passing (MMP) layers. In short, MMP layers invoke a series of standard message passing layers on a hierarchy of graphs corresponding to different characteristic lengthscales. This section briefly outlines the formulation of these layers. Before proceeding, it should be noted that the general multi-scale message passing concept, inspired by methods used in multigrid algorithms for solving partial differential equations, has been successfully applied in several mesh-based modeling works [16, 15, 34, 47].
@@ Appendix D Architecture and Training Details
Figure 16: Multiscale message passing (MMP) layer used in baseline and fine-tuned GNN architectures in Fig. 1.

### v4

@@ 3.2.3 Optimal Selection of Regularization Factor
The discussion above reveals the key role played by the regularization factor $\lambda$ as a controller for the trade-off between error tagging capability (quantitatively measured by the mean-squared error budget $\cal B$) and predictive accuracy of the interpretable GNN. The presence of such a trade-off – a direct by-product of the form of the regularized objective function in Eq. 5 – naturally leads to the notion of whether or not an ideal (or optimal) $\lambda$ that balances accuracy with error-tagging capability can be determined both a-priori and in a problem-independent manner. In particular, the following question comes into play: for a given Top-K reduction factor, is it possible to determine a value for $\lambda$ that maximizes the error budget contained in masked field (i.e., maximizes error-tagging capability) while also retaining the predictive capability of the baseline GNN? A $\lambda$-selection strategy to this end is described below.
To guide selection of $\lambda$, the expression in Eq. 5 can be simplified under two limiting conditions, both of which represent idealized performance targets of the regularized interpretable model. The first condition coincides with maximizing the MSE budget: in an ideal error-tagging scenario, 100% of the error in the forecast should be contained in the nodes identified by the masked field (a budget $\cal B$ of 1). In the ${\cal B}\rightarrow 1$ limit, the budget regularization loss ${\cal L}_{\cal B}$ also becomes 1. The second condition operates under the assumption of zero error in the interpretable model; in other words, an ideal model in the predictive sense would also recover ${\cal L}_{R}=0$ (see Eq. 5). Under these two limiting conditions, a simplified expression for the modeling objective in Eq. 5 can be obtained as $0={\cal L}_{\text{MSE}}+\lambda$, which in-turn provides an estimation for an ”ideal” $\lambda$ as $\lambda=-L_{\text{MSE}}$. Since a baseline model is available, the value $L_{\text{MSE}}$ can be set to $L_{\text{BL}}$, the converged MSE loss of the trained baseline GNN.
Although a negative value for $\lambda$ ($\lambda=-L_{\text{BL}}$) appears counter-intuitive, the relationship stems directly from the ideal modeling limits described above. Since $\lambda$ is negative, the budget regularization term during interpretability enhancement is adjusted to the direct budget instead of the inverse budget (i.e., ${\cal L}_{B}={\langle{\cal B}\rangle}$ instead of ${\cal L}_{B}=\frac{1}{\langle{\cal B}\rangle}$) to penalize large negative values for $\lambda{\cal L}_{B}$ during training. Although not shown here, the training loss histories for the MSE contribution (${\cal L}_{\text{MSE}}$, the single-step forecast error) validate this approach, in that the corresponding interpretable model in the $\lambda=-L_{\text{BL}}$ setting was found match the baseline prediction error on the training set.
@@ 3.3 Masked Fields on Extrapolated Geometries
The results in Sec. 3.1 show how the interpretability enhancement procedure provides the GNN the ability to generate identifiable structures through the masked fields. The results in Sec. 3.2 build on this intuition: through a budget regularization procedure conditioned on the identified masked fields, it was shown how a baseline GNN can be explicitly enhanced for interpretable error-tagging. Although these capabilities have many useful modeling implications (alongside the direct feature extraction benefits of the masked fields), trends in the previous sections were assessed only from the perspective of a single geometry, namely the backward-facing step.
@@
The objective of this section is to expand on the previous analyses by means of geometry extrapolation, which serves to not only verify masked fields trends, but also highlights the principal advantage of fully graph-based architectures: their ability to be used on unseen meshes. Here, interpretable GNNs trained on the BFS geometry are evaluated on two additional unseen geometries exhibiting different flow separation physics. These new geometries are the angled ramp and mounted cube configurations, which constitute a more challenging extrapolation task than, for example, extrapolating on another BFS geometry with the step positioned in a different part of the domain (the latter does not change the mechanism of flow separation, but rather only its spatial location). Fig. 14 in Appendix A provides images of the meshes used for the new geometries. Within the ramp and cube extrapolation context, the goal is to assess consistency in the masked fields from two angles, described below.
@@
Feature Identification and Error Tagging: To highlight error correlations, Fig. 11 provides instantaneous masked fields and corresponding spatial distributions of velocity component errors for the three aforementioned geometries: the backward-facing step, the ramp, and the wall-mounted cube. Analogous to Fig. 10, the figure includes outputs from interpretable GNNs trained with different budget regularization factors on the BFS geometry.
@@
For the BFS configuration in Fig. 11(a), the interpretable GNNs identify coherent structures downstream of the step. A similar identification of coherent structures is seen for the new geometries, in similar parts of the domain relative to the geometric feature. Most notably, the shift in the masked field towards the step anchor with increasing $\lambda$ – observed originally in BFS configuration – is also observed in the ramp configuration in Fig. 11(b). This shift is not seen in the cube case, which shows a markedly lower sensitivity to $\lambda$ than the other two geometries. Despite this, the GNN appears to be identifying coherent structures even in the extrapolated geometries, with the spatial location of these structures generally correlating with high levels of forecasting error. Additionally, the same concentration effect provided by the budget regularization strategy (which is the increase in the amount of error contained in the masked fields with $\lambda$) observed in the BFS case is observed in the ramp geometry – this effect is again much less apparent in the cube geometry. Although the masked fields are still picking up features in the mounted cube configuration, the lack of $\lambda$-sensitivity in Fig. 11(c) is likely due to the high absolute errors observed in the respective forecasts. Unlike the BFS and ramp cases, the cube-induced flowfield is not only governed by separation, but also the physical effects of flow impingement upstream of the geometry. Additionally, as described in Appendix A, the range in cells resolutions observed in the mesh for the cube configuration is greater than the other two cases, which may also be contributing to the higher errors. Ultimately, however, the extrapolations show how the masked fields largely preserve their feature identification capability in unseen geometries and largely remain correlated with regions of spatially concentrated forecasting error.
@@
Mask Decoherence During Rollout: Alongside analysis of spatial distributions and error tagging capability, it is of particular interest to evaluate whether the trends in the interpretable GNNs during rollout also translate to extrapolated geometries. More specifically, recall that through Fig. 7(d), presence of high-frequency noisy artifacts in the masked fields were shown to be intrinsically tied to the emergence of rollout errors in the BFS configuration. This emergence of noise is referred to as ”decoherency”. If a similar decoherency is observed during rollout for different geometries, the relationship between such a coherency breakdown and the presence of prediction errors is further verified (i.e., it strengthens confidence in the utility the masked fields provide in terms of predicting the error itself).
@@ 4 Conclusion
This work introduces an interpretability enhancement strategy for graph neural network based surrogate models, with focus on unstructured grid based fluid dynamics. More specifically, given a baseline GNN, the strategy consists of (a) appending an interpretable module to the GNN architecture through a skip connection, and (b) an isolated training stage to optimize the parameters introduced by this appended module. The resultant interpretable GNN adds interpretability to the baseline GNN by means of a graph pooling layer that acts as an adaptive sub-graph sampler, where these sub-graphs are identified in accordance to the underlying objective function (which here is the mesh-based forecasting error). Through this pooling operation, it is shown how the GNN forward pass generates masked fields as a by-product, which isolates regions in physical space in which these sampled sub-graphs are active. The features contained in these masked fields allow one to interpret the relationship between the input flow field at the current time instant, and the model-derived forecast at a future time instant. Through visualization of the masked fields, the interpretable GNNs allow one to access regions in physical space intrinsically linked to the modeling task (forecasting), thereby making the model behavior directly accessible from the perspective of the objective function used during training. As a result, the structures identified by pooling operation serve as an accessible link between the forecast-based objective function, the baseline model architecture, and the underlying physics of the problem being modeled.
@@
Additionally, by adding a regularization term to the objective function during the interpretability enhancement procedure in the form of an error budget, it was shown how the mask fields can also be used to identify, during inference, nodes that describe a majority of the anticipated forecasting error. Since the masked fields are an explicit function of the input flowfield (and not the predicted future flowfield), this regularization step adds a form of interpretable error-tagging capability to the modeling framework. A strategy to guide optimal selection of the regularization hyperparameter was also introduced, resulting in a novel approach that adds error-tagging capability to baseline GNNs without sacrificing baseline prediction accuracy. Additional analysis was performed on unseen geometries, with the goal of verifying the identification and error-tagging capability of the masked fields in challenging extrapolation scenarios. The analysis revealed how key GNN properties – particularly error localization and mask field decoherency – can indeed translate to unseen geometries.
The authors believe that the strategies introduced here begin to address the limitations of existing black-box mesh-based modeling strategies, and can readily be extended to other applications both within and outside of fluid dynamics surrogate modeling. It should be emphasized that although the identification of coherent structures is enabled by the Top-K sub-sampling mechanism, such identification is ultimately tied down to the physics being modeled (represented by both the domain configuration through initial and boundary conditions, and the objective function used during optimization). Depending on the problem, it is up to the end-user to assign physical meaning to the identified structures. For this reason, the authors believe that a similar coherent identification property of the masked fields is expected to be preserved in other physical applications/configurations, particularly those characterized by dominant instantaneous features that contribute significantly to the dynamics of the system in question (e.g., shock waves, detonation waves, vortices, etc). For more complex turbulent flows in three spatial dimensions and at high Reynolds numbers, it is unclear if the method will produce masked fields that identify (a) many disjoint small patches in physical space, (b) a few disjoint large patches, or (c) some combination of the two; further investigation in this direction is required.
@@
To conclude, there are many avenues for future work. The objective functions studied here concentrated on single-step forecasting errors – an interesting direction for future work is to append the same interpretable sub-sampling module, but include more rollout steps during the training process, thereby producing an interpretable stability enhancement. Including physics-based loss terms is another promising direction, as it would allow one to investigate the relationship between the identified structures and governing equation residuals. Additionally, although the modeling context in this work focused on a node-level regression task, extension of the ideas presented here for graph-level (i.e., more macroscopic) regression tasks may be possible. Regarding treatment of multiple geometries, a promising direction is to enhance a GNN for interpretability on geometry B using a baseline trained on geometry A. Further, so long as the input is represented as a graph, the authors believe that the interpretability and error indication enhancements developed here can be readily extended to different graph-based (and more complex sequence-based) architectures, such as graph transformers. Lastly, from the feature identification angle, establishing a connection between the Top-K procedure used here and other unsupervised feature extraction approaches (based on data-driven balance models [46] and physics-informed clustering [47], for example) is an interesting direction. These aspects, as well as other directions, are left for future studies.
@@ 5 Acknowledgements
The manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (Argonne). The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable world-wide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. This work was supported by the U.S. Department of Energy (DOE), Office of Science under contract DE-AC02-06CH11357. This research used resources of the Argonne Leadership Computing Facility, which is a U.S. Department of Energy Office of Science User Facility operated under contract DE-AC02-06CH11357. RM acknowledges funding support from ASCR for DOE-FOA-2493 “Data-intensive scientific machine learning” and from the RAPIDS2 Scientific Discovery through Advanced Computing (SciDAC) program.
@@ Appendix B Data Generation with OpenFOAM
As described in the beginning of Sec. 2, using the OpenFOAM solver, training and testing data was generated in the step height Reynolds number range of [26214, 45589]. As a reference, for this step height Reynolds number range of [26214, 45589], the corresponding friction Reynolds number range was found to be [3296, 4986]. To compute the friction Reynolds number, a characteristic length scale of half the initial channel height was used (half the channel height before encountering the step). The characteristic velocity was computed by taking a space-time average of the friction velocity using quantities at the lower wall upstream of the step anchor; space averages were taken in the streamwise (x) coordinate range of [$-5L_{s}$, 00], where $L_{s}$ is the step height and 0 is the x-coordinate location of the step anchor. Friction velocities were computed with the help of the wallShearStress post-processing function in OpenFOAM.
@@ Appendix C Multiscale Message Passing Layer
As illustrated in Fig. 1, the baseline GNN architecture and the interpretable sub-sampling module leverage multiscale message passing (MMP) layers. In short, MMP layers invoke a series of standard message passing layers on a hierarchy of graphs corresponding to different characteristic lengthscales. This section briefly outlines the formulation of these layers. Before proceeding, it should be noted that the general multi-scale message passing concept, inspired by methods used in multigrid algorithms for solving partial differential equations, has been successfully applied in several mesh-based modeling works [16, 15, 35, 51].
@@ Appendix D Architecture and Training Details
Figure 16: Multiscale message passing (MMP) layer used in baseline and interpretable GNN architectures in Fig. 1.
@@
Parameter counts and computational costs: The baseline model (which includes node and edge encoders, node decoder, and 2 MMP layers – refer to Fig. 1) contains $2.41\times 10^{6}$ parameters. The interpretable sub-sampling module parameters are populated by a single Top-K layer and 1 MMP layer, producing $1.16\times 10^{6}$ parameters and a value of 0.48 for the ratio (# interpretable sub-sampling module parameters / # baseline model parameters), which is fixed across all experiments. It should be noted that in the interpretable module, practically all parameters stem from the MMP layer, with only $N_{F}=128$ (the hidden graph node feature dimensionality) parameters contained in the Top-K layer (the projection vector); as such, this ratio can be readily reduced by specifying a lower-complexity MMP layer in the interpretable sub-sampling module.
Training times will ultimately depend on hardware, batch sizes, and dataset/graph sizes. However, for reference, using 8 Nvidia A100 GPUs, a training set size of roughly 40k snapshots, and a graph batch size of 2 (per GPU), the training time for a single epoch using the baseline GNN was 106.4 seconds, and the corresponding training time for the interpretable GNN was 90.2 seconds. Although the model size itself is larger, the lower training time in the interpretable GNN comes from the fact that all baseline model parameters are frozen (rendering a less intensive back-propagation). On the other hand, in the inference stage, it was found that the cost of interpretable GNNs were roughly 2x the cost of the baseline (roughly the same proportion as the parameter ratio).
Appendix E Masked Field Sensitivity and Stability
The tanh activation function was used in the Top-K layer (refer to Fig. 2) as it is the recommended/default setting in the PyTorch Geometric implementation [52]. However, it is important to understand the sensitivity of the masked fields to the activation function employed during the pooling stage. To shed some light on this front, this section provides additional numerical experiments that replace the tanh activation in the Top-K layer with a relu activation, with the goal of understanding the effect of this change on the identified masked fields.
Figure 17 shows a comparison of masked fields generated from single-step predictions at two different time instances using these activation settings (to facilitate this comparison, interpretable GNNs were trained using at the same Top-K reduction factor of RF=16 without any budget regularization, i.e., $\lambda=0$). The figure highlights how the identified regions are indeed sensitive to the activation function used, in that the tanh model (used throughout the main text) picks up additional structures in the far-downstream region of the step that are not identified by the relu model. However, despite this, the key structures near the step and in the near-downstream region are picked up by both models, and are qualitatively similar.
It should be noted that, although not shown here, the GNN leveraging the relu activation in the Top-K layer resulted in a higher overall loss (worse performance) than the tanh counterpart, which is likely the major contributor to the differences in the masked fields in Fig. 17. A similar trend is observed in Fig. 10 in the main text: the models with higher values of $\lambda$, when compared to the $\lambda=0$ model, also produce a higher MSE loss and do not pick up the far-downstream structures.
Ultimately, Fig. 17 suggests that changing the activation function in the Top-K layer does not fundamentally alter the model’s ability to pick up key coherent structures in the masks, though the authors recommended using tanh in the Top-K layer for these applications. It is hypothesized that the bounded nature of the tanh function helps in the training process for the Top-K vector, but a more comprehensive study on the effect of activation functions is warranted, as with any neural network based model (there are many more candidate functions to choose from).
Figure 17: (Left) Masked field visualizations comparing interpretable GNN trained using tanh activation in the Top-K layer (top) and relu activation in the Top-K layer (bottom). (Right) Same as left, but for a different input velocity field (different snapshot). Both snapshots are extracted at Re=35392 (trends were found to be similar at all other tested Reynolds numbers).
The above discussion is centered around characterizing the stability of the masked field in response to architectural changes. An alternative viewpoint (that also serves to verify the physical significance of the identified structures) is to extract the masked field sensitivity to parameter perturbations in the same GNN architecture. To this end, Fig. 18 shows spatial distributions of the standard deviations of masked fields produced using an ensemble of 30 different interpretable GNN models of the same architecture (i.e., each member in the ensemble has the same architecture as used in the main text, with tanh activations in the Top-K layer). The model ensemble was constructed by first taking a converged interpretable GNN model (RF=16, $\lambda=0$), and then performing stochastic gradient descent (SGD) training iterations on the converged model with a small learning rate of $10^{-7}$. The model ensemble was then populated by saving the parameter states at each of these SGD iterations (this strategy has been used in previous work for uncertainty quantification of machine learning models [56]). The result is a set of models with the same effective performace/loss, but “perturbed” parameter states, offering a useful testbed for studying the stability/consistency of structures in the masked fields through statistical analysis.
Figure 18: Masked field standard deviations during rollout steps 1, 30, and 50 at Re=35392. Standard deviations produced using RF=16 interpretable GNNs with $\lambda=0$.
The plots in Fig. 18 are computed by feeding the same input to each of the models in the ensemble, collecting the respective masks (there will be 30 masks at each time instant, since there are 30 models in the ensemble), and then computing the node-wise standard deviation in the masks. Figure 18 shows these standard deviations at three different rollout steps (1, 30, and 50) using an initial condition from the Re=35392 trajectory. Interestingly, the figures imply that the masks produced by the Top-K sub-sampling are relatively stable until about 30 steps, with noticeable standard deviations appearing at 50 steps and beyond concentrated around the near-step region. The emergence of the mask standard deviation at around 25-30 steps correlates with the high rate of growth in the masked field decoherence shown in Fig. 12 in the original manuscript near roughly the same point. Ultimately, the results in Fig. 18 serve to verify stability in the masked field structures in the early rollout stages. Additionally, this initial assessment reveals a promising direction in the form of connecting masked field decoherence in a single model to ensemble-based model uncertainties. A more detailed exploration of this connection is left for future studies.